OpenStudy (anonymous):

What is the X for in this problem: 6 1/3 X + 1 5/6 = 7 2/3

OpenStudy (anonymous):

\[6 \frac{ 1 }{ 3 }x + 1\frac{ 5 }{ 6 }=7\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

et's solve your equation step-by-step. (6)(1) 3 x+ (1)(5) 6 = (7)(2) 3 Step 1: Simplify both sides of the equation. 2x+ 5 6 = 14 3 Step 2: Subtract 5/6 from both sides. 2x+ 5 6 − 5 6 = 14 3 − 5 6 2x= 23 6 Step 3: Divide both sides by 2. 2x 2 = 23 6 2 x= 23 12 Answer: x= 23 12

OpenStudy (anonymous):

Your answer is x=23/12

OpenStudy (anonymous):

So I divice it by the LCD?

OpenStudy (anonymous):

Divide*

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

Ill take a medal :)